Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 110 miles per hour. Train A passes a station at 4:20 A.M. If train B passes the same station at 4:50 A.M., at what time will train B catch up to train A?
Answer by mananth(16949)

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 110 miles per hour. Train A passes a station at 4:20 A.M. If train B passes the same station at 4:50 A.M., at what time will train B catch up to train A?
...
Train A 4.20 A.M speed = 100 mph
Train B 4.50 A.M speed = 110 mph
..
difference in time they reach station = 30 minutes = 1/2 hour.
...
catch up speed of train B will be 110-100 = 10 mph.
..
Train A has traveled 50 miles in 1/2 hour.
...
catch up time = 50/10 = 5 hours.
Train B will catch up at 9;50 A.M
...
m.ananth@hotmail.ca